# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

### v1

@@ Abstract
Results. We present the NEMESIS catalogue of YSOs for the OSFC, which includes data collated for 27,879 sources covering the whole mass spectrum and the various stages of pre-Main Sequence evolution from protostars to diskless young stars. The catalogue includes a large collection of panchromatic photometric data processed into spectral energy distributions, stellar parameters (, spectral types, , , RV), infrared classes, equivalent widths of emission lines related to YSOs accretion and star-disk interaction, and absorption lines such as lithium and lines related to source’s gravity, X-ray emission observables, photometric variability observables (e.g., variability periods and amplitudes), and multiplicity labels.
@@
As astronomy dives into the era of Big Data, we are now faced with the challenges of reliable applications of artificial intelligence and machine learning techniques. A major obstacle for unbiased machine learning applications in the domain of star formation stems from the data available to validate results, which are often susceptible to small numbers or limited to the specifications of a few larger-scale surveys. To address the scarcity of polyvalent databases of young stars, we revisited the published literature targeting the Orion star-formation complex (OSFC) under the framework of the NEMESIS to build a reference sample Young Stellar Objects (YSOs) for use in upcoming research. We hereby report the work for data curation and a general description of the NEMESIS catalogue of young stellar objects for the Orion Star Formation Complex.
@@
NEMESIS (New Evolutionary Model for Early stages of Stars with Intelligent Systems) is an H2020 project111https://nemesis.univie.ac.at/ aiming at revisiting the star formation paradigm with the aid of Big Data and machine learning techniques. NEMESIS efforts included a new version of the Herschel/PACS point source catalogue (Marton et al., 2024a) with an associated Deep Neural Network approach for source removal (Madarász et al., 2024 submitted); the study of the 2D morphology of YSOs in photometric images using self-organising maps (Hernandez et al., 2024 submitted), the variability characterisation of an all-sky sample YSOs from Gaia DR3 (Mas et al., 2025 in prep), a compendium of the application of deep learning techiniques for the identification of YSOs all sky (Marton et al., 2024b, 2025 in prep) with the three latter presenting applications of the present catalogue. Additionally, two follow-up studies added value to the present catalogue through the fitting of YSO’s spectral energy distributions (SEDs) to synthetic libraries based on radiative transfer models (Gezer et al., 2025 in prep), and the ranking of bona fide YSOs (Roquette et al. in prep. 2025).
The Orion Star Formation Complex (OSFC; Bally, 2008) was chosen as our target as it is the largest and most diverse nearby star formation region within 500 pc of the Sun, with tenths of thousands of YSOs distributed over almost 600 square degrees in the sky (see Fig. 1). The OSFC harbours populations that span all ages relevant to studying young stellar evolution ( Myr; Kounkel et al., 2018) and covers the whole stellar mass spectrum from high-mass stars to intermediate- and low-mass YSOs (Hillenbrand, 1997; Muench et al., 2008; O’Dell et al., 2008).
@@
Our data curation (Sec. 2) started from a historical data compilation based on peer-reviewed articles focused on studying YSOs in the OSFC (Sec. 2.1). This historical compilation was then complemented with data from large photometric and spectroscopic surveys (Sec. 2.2 and 2.3) and from photometric data available at the Centre de données astronomiques de Strasbourg (CDS; Genova et al., 2000, Sec. 2.4). The motivations and criteria for the inclusion of sources in our compilation are fully discussed in Sec. 3, along with the justification of the data types collated into our catalogue. These include stellar parameters, SEDs, equivalent widths (EW) of emission lines and other spectroscopically derived features relevant to the study of YSOs, infrared (IR) classes related to YSOs’ evolutionary stage and information on their X-ray emission - the collated data are also summarised in App. C. We further added value to our catalogue by employing the curated SEDs to homogeneously derive IR classes for of the sources in the catalogue (Sec. 3.3.1) and by utilising a multifaceted approach to evaluate and multiplicity of sources (Sec. 4). Finally, in Sec. 5 we discuss the incidence of massive stars and contamination by other types of sources in our catalogue.
@@ 2 Data Curation
We reviewed the published literature targeting OSFC with the goal of building the largest catalogue of candidates and confirmed YSOs in the region. Our definition of the OSFC field encompasses 564 square degrees that cover the box between RA 74.2 and 92 deg and DEC and deg. We considered YSOs explicitly listed as members of the Monoceros R2 region (at RA, DEC: 91.94825, -6.37850 deg) to be outside our scope, although some of those were still indirectly included in our compilation. The sky distribution of 27,879 collected YSOs is shown in Fig. 1.
@@ 2.1 Data mining strategy for historical data
At the high stellar-mass end, our compilation included both higher-mass YSOs, such as Herbig Ae/Be, as well as O-type stars, which albeit no longer young within their own evolutionary path, are a youth indicator for a coeval population covering the full mass spectrum. Whenever possible, we have identified and flagged massive stars inferred to be beyond their zero-age main sequence (ZAMS) (see also Sec. 5.2.1).
@@ 2.1.1 Identification of Relevant Datasets through the NASA/ADS Database
Next, we applied a “bag-of-words” method to process these bibliography entries. We started by merging titles, abstracts, and keywords for each bibliographic entry into a single text-based entry. We then employed Natural language processing (NLP) methods to tokenise, lemmatisate, merge multi-word terms into single tokens, remove stop words, and homogenise spelling. The processed text was then scored on the basis of the occurrence of contextual tokens provided in an input vocabulary. This input vocabulary included terms such as Young Stellar Object, Star Formation, Young, Equivalent Width, Emission Line, Line Profile, Pre Main Sequence, Accretion, H-alpha, T Tauri, Herbig, Disk, protoplanetary, IR-excess and protostar along with alternative versions of these terms, including their abbreviations, synonyms, and derived words.
@@ 2.1.3 Data Integration
We processed the data included in the historical compilation to build a master catalogue in which each YSO-candidate previously identified in the literature is attributed a unique identifier. Sec. 3 provides more details of this processing. For each new literature reference integrated into our compilation, the new data set was matched to the master catalogue preferentially based on the identifier codes defined in previous studies. When unique identifiers were unavailable, we employed TOPCAT’s Pair-Match to join tables based on coordinate matching in the FK5 system. We used a 2” matching radius as the default. After testing different radius to match publications sharing common identifier codes, this was most often the best one. However, in cases where the source publication explicitly described larger PSF or astrometric precision, a larger radius of up to 10” was required.
@@ 2.2 Data from large photometric surveys
In addition to the compilation of YSO candidates, we have also built a reference photometric database for the OSFC field based on a list of large-scale photometric surveys, including observations at 34 photometric bands and covering the range wavelength range 0.15–160 . This allowed us to build a regular baseline for our panchromatic compilation, helping us to understand our historical compilation’s detection limits and completeness. Furthermore, by retrieving data directly from these surveys’ archives and following their user documentation, we could guarantee that the catalogues were properly cleaned before they were integrated into our database. It also helped verify the reliability of the data retrieved using the VizieR-SED tool (Sec. 2.4). Unless stated otherwise, we used a 2” matching radius to identify counterparts in our catalogue.
@@ 2.3 Data from large spectroscopic surveys
Gaia DR3 includes spectroscopic data products based on low-resolution spectra from the Blue and Red Photometers (BP/RP spectra ; Andrae et al., 2023; De Angeli et al., 2023) and medium resolution spectra from the Radial Velocity Spectrometer (RVS; Recio-Blanco et al., 2023). Along with a series of stellar parameters, the former includes the emission line (Sec. 3.6) with pseudo-equivalent widths estimated as part of the Extended Stellar Parametrizer for Emission-Line Stars (ESP-ELS) (Creevey et al., 2023), and the latter includes the Ca II infrared triplet (Lanzafame et al., 2023).
@@ 2.4 VizieR Photometry Viewer
We have further complemented our database with photometric data collected using the VizieR Photometry viewer. Powered by CDS, VizieR-SED is a tool intended for the visualisation of photometry around a given sky position. Hence, for an input coordinate or source name, it returns photometric observations “extracted around a sky position from photometry-enabled catalogues in VizieR” (VizieR Photometry viewer, ), where this photometry data can be exported in typical flux density units. The interpretation of the tool’s output as an SED is cautioned as CDS provides “no guarantee that all photometry points correspond to the target, especially for extended sources” (VizieR Photometry viewer, )333As part of our experience with VizieR-SED we second this recommendation and strongly advise against the indiscriminate use of SEDs retrieved with the tool for scientific purposes that require precise SEDs. One of the issues we identified can be traced back to the zero points adopted for magnitude-flux conversions in the tools’ back-end. The zero point values used can be recovered from the VizieR meta data tables METAfltr and METAphot at VizieR’s ReferenceDirectory. For example, 2MASS zero points adopted by CDS can be traced back to Pickles & Depagne (2010), and are offset in relation to official zero point values recommended by the survey’s documentation (Cohen et al., 2003), yielding a offset in the flux densities converted by VizieR.. Nevertheless, VizieR-SED provides a powerful tool for data mining the CDS database, and we have employed it to further search for relevant archival photometry for our sources. For that, we have followed their API’s recommendation to develop a Python script to query VizieR around the coordinates of each of our YSO candidates. We initially used a search radius of 5”. This wider search radius was chosen purposely, as it is prone to including contaminants from neighbouring sources, and we have taken advantage of this to flag sources likely susceptible to such contamination (see Sec. 4.4).
@@
For the specific purpose of complementing the SEDs of sources in our catalogue (Sec. 3.2), we have reduced this radius to 2” to minimise contamination by neighbouring sources. We visually inspected a large number of SEDs produced with VizieR-SED data and compared them with SEDs built using photometry from our historical compilation (Sec. 2.1) and large photometric surveys (Sec. 2.2), which had their magnitude-flux conversions carefully carried out by ourselves using meta-data available in their original publication. This procedure allowed us to identify and exclude from the SEDs data from surveys showing large systematics compared to the larger body of data available. We have post-processed VizieR-SED outputs to remove data coming from all CDS tables already included in the compilation (i.e., from Sec. 2.1 and 2.2).
@@
We also noticed large spreads in flux density in the SEDs generated by VizieR-SED, especially in the optical range 0.4-0.8 m. By further inspecting these data original tables, we verified that although a fraction of the spread remained unexplained, most of these could be traced back to either multi epoch surveys (YSOs are typically photometric variables as discussed in Sec. 3.8) or to surveys targeting extended sources and publishing photometry extracted for different aperture sizes (e.g. SDSS). Finally, we opted to avoid this wavelength range and kept only data in the ranges 0.1-0.4 m and 0.8-1000m, which seemed less affected by these issues.
@@
We have also identified a significant number of duplicated data, introduced by scientific publications re-reporting photometric data from large surveys. For example, although we have removed all 2MASS official data release tables - as this data has already been included in our compilation in Sec. 2.2 - many sources had VizieR SEDs with dozens of 2MASS JHKs data points, most of which having the same flux values but originating from different CDS tables. Together with multi-epoch data, this type of duplicate is problematic for quantities derived from SED-fits, as in the added-value quantities discussed in Sec. 3.3.1, as it biases the fit results towards the filters with a larger number of photometric points. We have, thus, further processed the SEDs from VizieR-SED to aggregate photometric measurements reported under the same filter name by averaging their fluxes weighted by their uncertainties and propagating these uncertainties accordingly.
@@ 3 Criteria for inclusion in this compilation and description of historical data-types
Throughout our historical data curation (Sec. 2.1), we focused on sources in the OSFC field previously reported as young in the literature. We have adopted a positive evidence rule for inclusion in the catalogue, where a given source had to be reported as a YSO candidate by any but at least one of the types of study discussed in this section. Our approach thus prioritises completeness over purity. In this section, we summarise the main YSO identification approaches employed by the scientific publications included in our historical compilation. We note that a revision of the different criteria adopted in the literature was beyond our scope. However, we were still interested in collating data that could support the ranking sources as bona fide YSOs. Hence, in this section, we also discuss a series of data types collected for this purpose. The various data types included in the catalogue are also summarised in Apps. B and C.
@@ 3.1.1 Spectral Types
Our historical compilation yielded spectral types for 7,485 sources. With the complementation by large surveys in Sec. 2.3, a total of 11,497 sources have spectral types. We tabulated these spectral types in the MK system, with luminosity class, or peculiar spectral characteristics, provided as an extra flag when available. When multiple previous studies provided independent spectral type derivations, a list of values is reported.

### v2

@@ Abstract
Results. We present the NEMESIS catalogue of YSOs for the OSFC, which includes data collated for 27 879 sources covering the whole mass spectrum and the various stages of pre-Main Sequence evolution from protostars to discless young stars. The catalogue includes a large collection of panchromatic photometric data processed into spectral energy distributions, stellar parameters (, spectral types, , , RV), infrared classes, equivalent widths of emission lines related to YSOs accretion and star-disc interaction, and absorption lines such as lithium and lines related to source’s gravity, X-ray emission observables, photometric variability observables (e.g., variability periods and amplitudes), and multiplicity labels.
@@
As astronomy dives into the era of Big Data, we are now faced with the challenges of reliable applications of artificial intelligence and machine learning techniques. A major obstacle for unbiased machine learning applications in the domain of star formation stems from the data available to validate results, which are often susceptible to small numbers or limited to the specifications of a few larger-scale surveys. To address the scarcity of polyvalent databases of young stars, we revisited the published literature targeting the Orion star-formation complex (OSFC) under the framework of the NEMESIS to build a reference sample of Young Stellar Objects (YSOs) for use in upcoming research. We hereby report the work for data curation and a general description of the NEMESIS catalogue of young stellar objects for the Orion Star Formation Complex.
@@
NEMESIS (New Evolutionary Model for Early stages of Stars with Intelligent Systems) is an H2020 project111https://nemesis.konkoly.hu/ aiming at revisiting the star formation paradigm with the aid of Big Data and machine learning techniques. NEMESIS efforts included a new version of the Herschel/PACS point source catalogue (Marton et al., 2024a) with an associated Deep Neural Network approach for source removal (Madarász et al., 2025); the study of the 2D morphology of YSOs in photometric images using self-organising maps (Hernandez et al., 2024 submitted), the variability characterisation of an all-sky sample YSOs from Gaia DR3 (Mas et al., 2025), a compendium of the application of deep learning techniques for the identification of YSOs all sky (Marton et al., 2024b, 2025 in prep) with the three latter presenting applications of the present catalogue. Additionally, two follow-up studies added value to the present catalogue through the fitting of YSO’s spectral energy distributions (SEDs) to synthetic libraries based on radiative transfer models (Gezer et al., 2025), and the ranking of bona fide YSOs (Roquette et al. in prep. 2025).
The Orion Star Formation Complex (OSFC; Bally, 2008) was chosen as our target as it is the largest and most diverse nearby star formation region within 500 pc of the Sun, with tenths of thousands of YSOs distributed over almost 600 square degrees in the sky (see Fig. 1). The OSFC harbours populations that span all ages relevant to studying young stellar evolution ( Myr; Kounkel et al., 2018) and covers the whole stellar mass spectrum from high-mass stars to intermediate- and low-mass YSOs (e.g., Hillenbrand, 1997; Muench et al., 2008; O’Dell et al., 2008). Previous research reported relatively large samples of YSOs located in the OSFC, but these were susceptible to the methodological specificities and limitations inherent to certain instruments or focused on subregions. For example, studies carried out with the Spitzer space telescope (e.g., Megeath et al., 2012, 2016; Fang et al., 2013; Getman et al., 2018) collectively identified YSOs. However, these are limited by Spitzer’s reduced coverage of the OSFC and favour the detection of less evolved YSOs with significant contributions of a disc or envelope to their SEDs. Large-scale kinematic surveys based on Gaia (e.g. Zari et al., 2018) characterised a similar number of YSOs, but are conversely biased towards optically visible sources. Similarly, large-scale spectroscopic surveys (e.g., Da Rio et al., 2016; Kounkel et al., 2018; Briceño et al., 2019; Hernández et al., 2023) focussed on the brightest sources located in less crowded areas and are accordingly biased toward more evolved YSOs, which already emerged from their envelopes, being near-infrared bright and optically visible. With the NEMESIS Catalogue of YSOs for the OSFC, we propose to go beyond these limitations by compiling a cumulative source list from the ensemble of previous research and combining complementation from public datasets and modern data science techniques to and mitigate, as possible, the biases from previous surveys.
@@
Our data curation (Sect. 2) started from a historical data compilation based on peer-reviewed articles focused on studying YSOs in the OSFC (Sect. 2.1). This historical compilation was then complemented with data from large photometric and spectroscopic surveys (Sect. 2.2 and 2.3) and from photometric data available at the Centre de données astronomiques de Strasbourg (CDS; Genova et al., 2000, Sect. 2.4). The motivations and criteria for the inclusion of sources in our compilation are fully discussed in Sect. 3, along with the justification of the data types collated into our catalogue. These include stellar parameters, SEDs, equivalent widths (EW) of emission lines and other spectroscopically derived features relevant to the study of YSOs, infrared (IR) classes related to YSOs’ evolutionary stage and information on their X-ray emission - the collated data are also summarised in App. D. We further added value to our catalogue by employing the curated SEDs to homogeneously derive IR classes for of the sources in the catalogue (Sect. 3.3.1) and by utilising a multifaceted approach to evaluate and multiplicity of sources (Sect. 4). Finally, in Sect. 5 we discuss the incidence of massive stars and contamination by other types of sources in our catalogue.
@@ 2 Data Curation
We reviewed the published literature targeting OSFC with the goal of building the largest catalogue of candidates and confirmed YSOs in the region. Our definition of the OSFC field encompasses 564 square degrees that cover the box between RA 74.2 and 92 deg and DEC and deg. We considered YSOs explicitly listed as members of the Monoceros R2 region (at RA, DEC: 91.94825, -6.37850 deg) to be outside our scope, although some of those were still indirectly included in our compilation. The sky distribution of 27,879 collected YSOs is shown in Fig. 1. Fig. 2 presents a schematic representation of the data curation workflow further described in this section.
@@ 2.1 Data mining strategy for historical data
At the high stellar-mass end, our compilation included both higher-mass YSOs, such as Herbig Ae/Be, as well as O-type stars, which albeit no longer young within their own evolutionary path, are a youth indicator for a coeval population covering the full mass spectrum. Whenever possible, we have identified and flagged massive stars inferred to be beyond their zero-age main sequence (ZAMS) (see also Sect. 5.2.1).
@@ 2.1.1 Identification of Relevant Datasets through the NASA/ADS Database
Next, we applied a “bag-of-words” method to process these bibliography entries. We started by merging titles, abstracts, and keywords for each bibliographic entry into a single text-based entry. We then employed Natural language processing (NLP) methods to tokenise, lemmatisate, merge multi-word terms into single tokens, remove stop words, and homogenise spelling. The processed text was then scored on the basis of the occurrence of contextual tokens provided in an input vocabulary. This input vocabulary included terms such as Young Stellar Object, Star Formation, Young, Equivalent Width, Emission Line, Line Profile, Pre Main Sequence, Accretion, H-alpha, T Tauri, Herbig, disc, protoplanetary, IR-excess and protostar along with alternative versions of these terms, including their abbreviations, synonyms, and derived words.
@@ 2.1.3 Data Integration
We processed the data included in the historical compilation to build a master catalogue in which each YSO-candidate previously identified in the literature is attributed a unique identifier. Sect. 3 provides more details of this processing. For each new literature reference integrated into our compilation, the new data set was matched to the master catalogue preferentially based on the identifier codes defined in previous studies. When unique identifiers were unavailable, we employed TOPCAT’s Pair-Match to join tables based on coordinate matching in the FK5 system. We used a 2” matching radius as the default. After testing different radius to match publications sharing common identifier codes, this was most often the best one. However, in cases where the source publication explicitly described larger PSF or astrometric precision, a larger radius of up to 10” was required.
@@ 2.2 Data from large photometric surveys
In addition to the compilation of YSO candidates, we have also built a reference photometric database for the OSFC field based on a list of large-scale photometric surveys, including observations at 34 photometric bands and covering the range wavelength range 0.15–160 . This allowed us to build a regular baseline for our panchromatic compilation, helping us to understand our historical compilation’s detection limits and completeness. Furthermore, by retrieving data directly from these surveys’ archives and following their user documentation, we could guarantee that the catalogues were properly cleaned before they were integrated into our database. It also helped verify the reliability of the data retrieved using the VizieR-SED tool (Sect. 2.4). Unless stated otherwise, we used a 2” matching radius to identify counterparts in our catalogue.
@@ 2.3 Data from large spectroscopic surveys
Gaia DR3 includes spectroscopic data products based on low-resolution spectra from the Blue and Red Photometers (BP/RP spectra ; Andrae et al., 2023; De Angeli et al., 2023) and medium resolution spectra from the Radial Velocity Spectrometer (RVS; Recio-Blanco et al., 2023). Along with a series of stellar parameters, the former includes the emission line (Sect. 3.6) with pseudo-equivalent widths estimated as part of the Extended Stellar Parametrizer for Emission-Line Stars (ESP-ELS) (Creevey et al., 2023), and the latter includes the Ca II infrared triplet (Lanzafame et al., 2023).
@@ 2.4 VizieR Photometry Viewer
We have further complemented our database with photometric data collected using the VizieR Photometry viewer. Powered by CDS, VizieR-SED is a tool intended for the visualisation of photometry around a given sky position. Hence, for an input coordinate or source name, it returns photometric observations “extracted around a sky position from photometry-enabled catalogues in VizieR” (VizieR Photometry viewer, ), where this photometry data can be exported in typical flux density units. The interpretation of the tool’s output as an SED is cautioned as CDS provides “no guarantee that all photometry points correspond to the target, especially for extended sources” (VizieR Photometry viewer, )333As part of our experience with VizieR-SED we second this recommendation and strongly advise against the indiscriminate use of SEDs retrieved with the tool for scientific purposes that require precise SEDs. One of the issues we identified can be traced back to the zero points adopted for magnitude-flux conversions in the tools’ back-end. The zero point values used can be recovered from the VizieR meta data tables METAfltr and METAphot at VizieR’s ReferenceDirectory. For example, 2MASS zero points adopted by CDS can be traced back to Pickles & Depagne (2010), and are offset in relation to official zero point values recommended by the survey’s documentation (Cohen et al., 2003), yielding a offset in the flux densities converted by VizieR.. Nevertheless, VizieR-SED provides a powerful tool for data mining the CDS database, and we have employed it to further search for relevant archival photometry for our sources. For that, we have followed their API’s recommendation to develop a Python script to query VizieR around the coordinates of each of our YSO candidates. We initially used a search radius of 5”. This wider search radius was chosen purposely, as it is prone to including contaminants from neighbouring sources, and we have taken advantage of this to flag sources likely susceptible to such contamination (see Sect. 4.3).
@@
For the specific purpose of complementing the SEDs of sources in our catalogue (Sect. 3.2), we have reduced this radius to 2” to minimise contamination by neighbouring sources. We visually inspected a large number of SEDs produced with VizieR-SED data and compared them with SEDs built using photometry from our historical compilation (Sect. 2.1) and large photometric surveys (Sect. 2.2), which had their magnitude-flux conversions carefully carried out by ourselves using meta-data available in their original publication. This procedure allowed us to identify and exclude from the SEDs data from surveys showing large systematics compared to the larger body of data available. We have post-processed VizieR-SED outputs to remove data coming from all CDS tables already included in the compilation (i.e., from Sect. 2.1 and 2.2).
@@
We also noticed large spreads in flux density in the SEDs generated by VizieR-SED, especially in the optical range 0.4-0.8 m. By further inspecting these data original tables, we verified that although a fraction of the spread remained unexplained, most of these could be traced back to either multi epoch surveys (YSOs are typically photometric variables as discussed in Sect. 3.8) or to surveys targeting extended sources and publishing photometry extracted for different aperture sizes (e.g. SDSS). Finally, we opted to avoid this wavelength range and kept only data in the ranges 0.1-0.4 m and 0.8-1000m, which seemed less affected by these issues.
@@
We have also identified a significant number of duplicated data, introduced by scientific publications re-reporting photometric data from large surveys. For example, although we have removed all 2MASS official data release tables - as this data has already been included in our compilation in Sect. 2.2 - many sources had VizieR SEDs with dozens of 2MASS JHKs data points, most of which having the same flux values but originating from different CDS tables. Together with multi-epoch data, this type of duplicate is problematic for quantities derived from SED-fits, as in the added-value quantities discussed in Sect. 3.3.1, as it biases the fit results towards the filters with a larger number of photometric points. We have, thus, further processed the SEDs from VizieR-SED to aggregate photometric measurements reported under the same filter name by averaging their fluxes weighted by their uncertainties and propagating these uncertainties accordingly.
@@ 3 Criteria for inclusion in this compilation and description of historical data-types
Throughout our historical data curation (Sect. 2.1), we focused on sources in the OSFC field previously reported as young in the literature. We have adopted a positive evidence rule for inclusion in the catalogue, where a given source had to be reported as a YSO candidate by any but at least one of the types of study discussed in this section. Our approach thus prioritises completeness over purity. In this section, we summarise the main YSO identification approaches employed by the scientific publications included in our historical compilation. We note that a revision of the different criteria adopted in the literature was beyond our scope. However, we were still interested in collating data that could support the ranking sources as bona fide YSOs. Hence, in this section, we also discuss a series of data types collected for this purpose. The various data types included in the catalogue are also summarised in Apps. C and D.
@@ 3.1.1 Spectral Types
Our historical compilation yielded spectral types for 7,485 sources. With the complementation by large surveys in Sect. 2.3, a total of 11,497 sources have spectral types. We tabulated these spectral types in the MK system, with luminosity class, or peculiar spectral characteristics, provided as an extra flag when available. When multiple previous studies provided independent spectral type derivations, a list of values is reported.